OpenStudy (anonymous):

find the equation of the line containing the two given points in standard form and slope- intercept form (-7,-4) and (14,2)

OpenStudy (anonymous):

For Standard Form Use: \[\large \frac{y - y_1}{y_2 - y_1} = \frac{x-x_1}{x_2 - x_1}\] Plug in the values: \((x_1, y_1) = (-7,-4)\) \((x_2, y_2) = (14,2)\)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

And then find slope: \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\] Then use: \[\large y-y_1 = m(x-x_1)\]

OpenStudy (anonymous):

ooh kk

OpenStudy (anonymous):

thanks you